OpenStudy (anonymous):

i need help answering this 3*(sq rt of 48) + 2*(sq rt of 12)

OpenStudy (anonymous):

\[16\sqrt{3}\]

OpenStudy (anonymous):

how do u solve

OpenStudy (anonymous):

\[3\sqrt{48}+2\sqrt{12}\] \[12\sqrt{3}+4\sqrt{3}\] \[16\sqrt{3}\]

OpenStudy (anonymous):

You need to factor out the perfect squares 48 is 3 * 16 which the 16 is 4 * 4 and 12 is 4 * 3 so you can factor the 4 out as 2 * 2 which leaves you with what sritama said.

OpenStudy (anonymous):

u explain it better than the book does

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

y is is 12(sq rt of 3) instead of 16(sq rt of 3) ????